can someone help with this basic econ questions:

1)Assume that you are planning on purchasing a new car for $30,000. You will be able to make a down payment of $3,300. The remaining $26,700 will be financed by the dealer. The dealer computes your monthly payment to be $615.00 for 48 months of financing. What is the dealer’s monthly rate of return on this car loan? What is the annual rate of return?

2)

A plant engineer is considering two machines to purchase:

Item Machine A Machine B
Initial Cost     $12,600 $17,350
Annual Savings     $5,500    $6,000
Annual Maintenance     $3,200 $2,500
Expected Life      7 years 7 years
Salvage Value     $1,750 $2,100

The firm's MARR is 11%. On the basis of the IRR criterion, which machine is the better choice?
Answer #1

Loan = 26700

Monthly payment = 615

Months = 48

Let month interest rate be i%, then

PW = - 26700 + 615*(P/A,i%,48) = 0

(P/A,i%,48) = 26700 / 615 = 43.414634

using trail and error method

When i = 0.25%, (P/A,i%,48) = 45.178695

When i = 0.5%, (P/A,i%,48) = 42.580318

using interpolation

i = 0.25% + (45.178695 - 43.414634) / (45.178695 - 42.580318) *(0.5 - 0.25)

i = 0.25% + 0.169727

i = 0.419% = 0.42% (Approx)

Nominal interest per year = 0.42 * 12 = 5.04% (Approx)

Pls upload the next question separately as per Chegg policy